# On The Structure of EFX Orientations on Graphs

### v1

@@ Abstract.
Fair division is the problem of allocating a set of items among agents in a fair manner. One of the most sought-after fairness notions is envy-freeness (EF), requiring that no agent envies another’s allocation. When items are indivisible, it ceases to exist, and envy-freeness up to any good (EFX) emerged as one of its strongest relaxations. The existence of EFX allocations is arguably the biggest open question within fair division. Recently, Christodoulou, Fiat, Koutsoupias, and Sgouritsa (EC 2023) introduced showed that EFX allocations exist for the case of graphical valuations where an instance is represented by a graph: nodes are agents, edges are goods, and each agent values only her incident edges. On the other hand, they showed NP-hardness for checking the existence of EFX orientation where every edge is allocated to one of its incident vertices, and asked for a characterization of graphs that exhibit EFX orientation regardless of the assigned valuations.
In this paper, we make significant progress toward answering their question. We introduce the notion of strongly EFX orientable graphs – graphs that have EFX orientations regardless of how much agents value the edges. We show a surprising connection between this property and the chromatic number of the graph, namely $\chi(G)$ for graph $G$. In particular, we show that graphs with $\chi(G)\leq 2$ are strongly EFX orientable, and those with $\chi(G)>3$ are not strongly EFX orientable. We provide examples of strongly EFX orientable and non-strongly EFX orientable graphs of $\chi(G)=3$ to prove tightness. Finally, we give a complete characterization of strong EFX orientability when restricted to binary valuations.
@@ 1. Introduction
Fair division is a problem of allocating a set $M$ of goods among $n$ agents in a fair manner [Steinhaus48]. This is an age-old problem with numerous contemporary applications, e.g., division of family inheritance [PrattZ90], divorce settlements [BramsT96], spectrum allocation [EtkinPT05], air traffic management [Vossen02], course allocation [BudishC10] and many more. [^1] See www.spliddit.org and www.fairoutcomes.com for a detailed discussion on fair division protocols used in day-to-day life.. In this paper, we study the fair division of indivisible goods. In this case, preferences of each agent $i\in[n]=\{1,2...n\}$ is represented by a monotone increasing valuation function $f_{i}:2^{M}\to\mathbb{R}_{\geq 0}$, specifying how much they value each subset of goods. An allocation of goods is a partition of $M$ into $n$ subsets $X=(X_{1},X_{2}...X_{n})$, where $X_{i}$ is allocated to agent $i$.
One of the most sought-after fairness notions is of envy freeness. An allocation is said to be envy free (EF) if no agent $i$ envies another’s allocation: $f_{i}(X_{i})\geq f_{j}(X_{j}),\forall j\in[n]$. With indivisible items, EF allocation may not exist, for example, allocating one iPhone among two agents who both value it highly. Hence, several relaxations of envy-freeness have been studied, arguably the strongest of which is envy-freeness up to any good (EFX). [Caragiannis19]. An allocation is said to be EFX if for any two agents $i,j$, $i$ does not envy any proper subset of $j$’s bundle. That is, for any $X\subset X_{j}$, $f_{i}(X_{i})\geq f_{i}(X)$. In other words, $i$ does not envy $X_{j}$ after removal of any good from $X_{j}$.
@@
An orientation of edges in $G$ can be interpreted as an allocation where each edge/good is given to its incident vertex/agent the edge is oriented towards. It is called an EFX orientation if the corresponding allocation is EFX. [Christodoulou2023] showed that an EFX orientation may not always exist and, furthermore, proved that determining whether a graph has an EFX orientation for a given valuation is NP-hard.
On the other hand, they showed that if we allow edges to be allocated to non-incident vertices, then an EFX allocation exist. However, they noted that EFX orientations are more desirable since they avoid lossy assignments (assigning an edge to a non-incident vertex). Motivated by this, they asked the following question [Christodoulou2023]:
@@ 1.1. Our Contributions
  • There exist graphs of with $\chi(G)=3$ which are not strongly EFX-orientable, as well as graphs with $\chi(G)=3$ which are strongly EFX-orientable, so this bound is sharp.
@@ 1.2. Further Related Work
Organization. In Section 2 we formally define the fair division problem and strong EFX-orientability. Section 3 discusses the necessary conditions for strong EFX-orientability and obtains the upper bound of 3 on the chromatic number. In the process it completely characterizes 0-1 strong EFX-orientability. Section 4 discusses the sufficiency conditions through bipartiteness or near-bipartiteness of the graph. Section 5 discusses structures in certain 3-chromatic graphs that prevent strong EFX-orientability, leaving 3-chromatic graphs as an undecidable case. Finally, we conclude with a brief discussion in Section 6.
@@ 2. Preliminaries
A discrete fair division instance is given by $([n],M,\mathcal{F})$, where $[n]$ is the set of agents, $M$ is the set of goods, and =$\mathcal{F}=\{f_{1},f_{2}..f_{n}\}$ is the set of valuation functions, one for each agent. For each $i\in[n]$, $f_{i}:2^{M}\rightarrow\mathbb{R}_{\geq 0}$ represents agent $i$’s value over the bundles (sets) of goods. Thus, $f_{i}$ is non-negative and monotone, i.e., for $A,B\subseteq M$, $A\subseteq B$ implies $0\leq f_{i}(A)\leq f_{i}(B)$. A valuation function $f_{i}$ is said to be additive if for every subset $X=\{m_{1},m_{2},...m_{k}\}\subseteq M$, $f_{i}(X)=f_{i}(\{m_{1}\})+f_{i}(\{m_{2}\})+...+f_{i}(\{m_{k}\})$, and a valuation is 0-1 additive if it is additive and $f_{i}(\{m\})\in\{0,1\}$ for all $m\in M$.
An allocation $X=(X_{1},X_{2}...X_{n})$ is a partition of $M$ where agent $i$ is assigned the bundle $X_{i}$. An allocation is said to be envy-free (EF), if $f_{i}(X_{i})\geq f_{i}(X_{j})$ for all $i,j\in[n]$. Since such an allocation may not exist when goods are discrete, we consider the following relaxation:
@@ 3.1. Characterization of 0-1 strongly EFX-Orientable Graphs
A graph $G$ is 0-1 strongly EFX-orientable if and only if, for every subgraph $H\subseteq G$ such that $H$ is a forest consisting of trees $T_{1},T_{2}...T_{k}$, for every $1\leq i\leq k$ there exists $x_{i}\in T_{i}$ such that $\bigcup_{i=1}^{k}N_{H}(x_{i})$ forms an independent set on $G$ (Figure 1 demonstrates this condition for a $K_{2,4}$ graph).
@@
($\Rightarrow$) For all the edges with asymmetric valuation, orient them towards the vertex that values the edge. Call the vertices that receive the asymmetrically valued edges special vertices. Let $H^{\prime}$ be the subgraph consisting of edges of weight 1 for both endpoints. For each component of $H^{\prime}$ that contains a special vertex, create a spanning tree with the special vertex as the root, and give each vertex the edge from its parent. For a component $C\subseteq H^{\prime}$ that contains a cycle but has no special vertex, we remove one edge $uv\in C$ where $uv$ is in a cycle, and construct a spanning tree through $C-uv$ with $v$ as the root. Assign each vertex in $C$ the edge from its parent, and assign $uv$ to $v$. Orient the rest of the edges in these components arbitrarily. Since all vertices in such components receive an edge they value, they do not envy anyone else, and are hence not envied by any vertex.
@@
($\Leftarrow$) Take a forest $H$ such that for its trees $T_{1},T_{2}...T_{k}$, for every collection of $x_{1},x_{2}...x_{k}$ such that $x_{i}\in T_{i}$, their neighborhoods in $H$ do not form an independent set on $G$. Set all the edges in $H$ to have weight 1 for both vertices, and set the rest of the edges to have weight 0 for both vertices. Note that any orientation on a tree must have a source, so for any orientation in $G$, we can choose a collection of sources in $s_{1},s_{2}...s_{k}$ in $H$ such that $s_{i}\in T_{i}$ for $1\leq i\leq k$. Since each $s_{i}$ receives nothing of value and has all its valued edges go to its neighbors, all vertices in $N_{H}(s_{i})$ are envied by $s_{i}$, so all vertices in $\bigcup_{i=1}^{k}N_{H}(s_{i})$ must be envied. However, since $\bigcup_{i=1}^{k}N_{H}(s_{i})$ does not form an independent set on $G$, then an edge of weight 0 must be between two vertices in that set. Assigning that edge will break EFX since both its endpoints are envied due to the orientation on $H$, hence such a graph is not 0-1 strongly EFX-orientable. ∎
@@
Let $G$ be a 0-1 strongly EFX orientable graph. For any matching $M$ on $G$, the subgraph induced by the vertices of $M$ have an independent set of size $|M|$.
@@
Let $M$ consist of edges $e_{1},e_{2},....e_{k}$. Each edge in a matching is a tree, and the edges form a forest. Hence, by Lemma 3.1, for each edge $e_{i}$ in $M$, we can select $v_{i}\in e_{i}$ such that $\bigcup_{i=1}^{k}N_{M}(v_{i})$ is an independent set. Since each vertex in $M$ has exactly one other neighbor also in $M$, then $\bigcup_{i=1}^{k}N_{M}(v_{i})$ consists of $|M|$ vertices, and hence is an independent set of size $|M|$. ∎
@@
Let $G$ be a graph that violates the matching condition in Corollary 3.2. Let $G^{\prime}$ the graph obtained by subdividing any edge of $G$ twice. Then $G^{\prime}$ violates the matching condition of Corollary 3.2 as well.
@@
Let the subdivided in edge on $G^{\prime}$ consist of $u,x_{1},x_{2},v$, where $uv$ is the original edge in $G$ and $x_{1},x_{2}$ are the added vertices. Suppose that for any matching in $G^{\prime}$, there exists an independent set of the same size on the graph induced by the vertices of that matching. Take a matching $M$ on $G$. We split this problem into two cases:
@@
We claim that $K_{4}$ violates the matching condition in Corollary 3.2. Indeed, the independence number of a $K_{4}$ is 1, while it’s possible to have a matching of size 2 in a $K_{4}$. By Lemma 3.3, a totally odd subdivision of a graph that violates the matching condition will violate the matching condition itself as well. Hence, 0-1 strongly EFX-orientable graphs cannot contain a totally odd $K_{4}$ subdivision, so $\chi(G)\leq 3$ by Lemma 3.4. ∎
@@ 4. Strong EFX-Orientability: Sufficient Condition via Bipartiteness
Note that $B$ must be an independent set, since $uv$ has one end in $A$, and all other edges go from $A$ to $B$, so no envy exists between vertices in $B$. No vertex in $A$ envies a vertex in $B$, since they got to take an edge before $B$, and the remaining edges incident to a vertex in $A$ were assigned to different neighbors of that vertex. While there might be envy towards vertices in $A$, every vertex in $A$ was assigned only one edge, making this orientation EFX. ∎
@@ 5. General Valuations: Forbidden Structures
Finally, the graph of two triangles connected by a path of length has a forest selected that consists of three trees, $T_{1}=v_{1}v_{3}v_{2}$, $T_{2}=v_{4}v_{5}$, and $T_{3}=v_{6}v_{7}$. Similar to the previous case, to select $x_{1},x_{2},x_{3}$ such that their neighborhoods in the forest form an independent set, $x_{1}$ cannot equal $v_{3}$ and must be either $v_{1}$ or $v_{2}$. This means that $x_{2}$ cannot equal $v_{5}$, since $v_{3}$ and $v_{5}$ are adjacent, so $x_{2}=v_{4}$. But both $v_{6}$ and $v_{7}$ are adjacent to $v_{5}$, so selecting $x_{3}$ is impossible, meaning that this graph is not 0-1 strongly EFX orientable as well. ∎
@@
Let the first triangle be $v_{1}v_{2}v_{3}$, and the second triangle be $v_{2}v_{3}v_{4}$. Let $v_{2}v_{3}$ have weight 0.5 for both endpoints, $v_{1}v_{2}$ and $v_{3}v_{4}$ have weight 1 for both endpoints, and the rest of the edges have weight 0 for both endpoints. All valuations are additive. Due to the symmetry of the graph, without loss of generality, we can assume that $v_{2}v_{3}$ is directed towards $v_{2}$. For the orientation to be EFX, $v_{2}$ cannot receive $v_{1}v_{2}$, as $v_{1}$ will envy $v_{2}$ even though $v_{2}$ has already received $v_{2}v_{3}$. Hence $v_{1}$ must receive $v_{1}v_{2}$, but $v_{1}$ will be envied by $v_{2}$, so $v_{1}v_{3}$ must be directed towards $v_{3}$. Since $v_{3}$ already has an edge, for similar reasons, $v_{3}$ cannot receive $v_{3}v_{4}$, so $v_{3}v_{4}$ is directed towards $v_{4}$. Note that $v_{2}$ and $v_{4}$ are both envied by $v_{2}$, but there is another edge between $v_{2}$ and $v_{4}$ that needs to be assigned, making an EFX-orientation impossible.
@@
Note while subdividing $e$, we created edges $ux_{1}$ and $vx_{2}$ which also have zero value, meaning that we can repeat this process again by subdividing any one of those edges twice, hence replacing $e$ with an odd path of edges alternating between zero and one values. The new graph will not have an EFX orientation with this valuation, and hence is not strongly EFX-orientable, as deisred. ∎
@@
Let $G$ be a $2-$connected bipartite graph with $|V(G)|\geq 4$. Suppose that an edge $e$ is added, which connects two vertices of the same partite set. Then $G+e$ is not strongly EFX-orientable.

### v2

@@ Abstract.
Fair division is the problem of allocating a set of items among agents in a fair manner. One of the most sought-after fairness notions is envy-freeness (EF), requiring that no agent envies another’s allocation. When items are indivisible, it ceases to exist, and envy-freeness up to any good (EFX) emerged as one of its strongest relaxations. The existence of EFX allocations is arguably the biggest open question within fair division. Recently, Christodoulou, Fiat, Koutsoupias, and Sgouritsa (EC 2023) showed that EFX allocations exist for the case of graphical valuations where an instance is represented by a graph: nodes are agents, edges are goods, and each agent values only her incident edges. On the other hand, they showed NP-hardness for checking the existence of EFX orientation where every edge is allocated to one of its incident vertices, and asked for a characterization of graphs that exhibit EFX orientation regardless of the assigned valuations.
In this paper, we make significant progress toward answering their question. We introduce the notion of strongly EFX orientable graphs – graphs that have EFX orientations regardless of how much agents value the edges. We show a surprising connection between this property and the chromatic number $\chi(G)$ of the graph $G$. In particular, we show that graphs with $\chi(G)\leq 2$ are strongly EFX orientable, and those with $\chi(G)>3$ are not strongly EFX orientable. We provide examples of strongly EFX orientable and non-strongly EFX orientable graphs of $\chi(G)=3$ to prove tightness. Finally, we give a complete characterization of strong EFX orientability when restricted to binary valuations.
@@ 1. Introduction
Fair division is a problem of allocating a set $M$ of goods among $n$ agents in a fair manner [Steinhaus48]. This is an age-old problem with numerous contemporary applications, e.g., division of family inheritance [PrattZ90], divorce settlements [BramsT96], spectrum allocation [EtkinPT05], air traffic management [Vossen02], course allocation [BudishC10] and many more  [^1] See www.spliddit.org and www.fairoutcomes.com for a detailed discussion on fair division protocols used in day-to-day life.. In this paper, we study the fair division of indivisible goods. In this case, the preferences of each agent $i\in[n]=\{1,2...n\}$ are represented by a monotone increasing valuation function $f_{i}:2^{M}\to\mathbb{R}_{\geq 0}$, specifying how much they value each subset of goods  [^2] Using $v_{i}$ to denote the valuation function is the convention in the literature, however, we will use $f_{i}$ to avoid confusing notation, as $v$ is also used to denote a vertex in a graph.. An allocation of goods is a partition of $M$ into $n$ subsets $X=(X_{1},X_{2}...X_{n})$, where $X_{i}$ is allocated to agent $i$.
One of the most sought-after fairness notions is of envy freeness. An allocation is said to be envy free (EF) if no agent $i$ envies another’s allocation: $f_{i}(X_{i})\geq f_{i}(X_{j}),\forall i,j\in[n]$. With indivisible items, an envy-free allocation may not always exist, for example, allocating one iPhone among two agents who both value it highly. Hence, several relaxations of envy-freeness have been studied, arguably the strongest of which is envy-freeness up to any good (EFX) [Caragiannis19]. An allocation is said to be EFX if for any two agents $i,j$, $i$ does not envy any proper subset of $j$’s bundle. That is, for any $X\subset X_{j}$, $f_{i}(X_{i})\geq f_{i}(X)$. In other words, $i$ does not envy $X_{j}$ after the removal of any good from $X_{j}$.
@@
An orientation of edges in $G$ can be interpreted as an allocation where each edge/good is given to its incident vertex/agent the edge is oriented towards. It is called an EFX orientation if the corresponding allocation is EFX. [Christodoulou2023] showed that an EFX orientation may not always exist and, furthermore, proved that determining whether a graph has an EFX orientation for a given set of valuations is NP-hard.
On the other hand, they showed that if we allow edges to be allocated to non-incident vertices, then an EFX allocation exists. However, they noted that EFX orientations are more desirable since they avoid lossy assignments (assigning an edge to a non-incident vertex). Motivated by this, they asked the following question [Christodoulou2023]:
@@ 1.1. Our Contributions
  • There exist graphs of $\chi(G)=3$ which are not strongly EFX-orientable, as well as graphs with $\chi(G)=3$ which are strongly EFX-orientable, so this bound is sharp.
@@ 1.2. Further Related Work
Organization. In Section 2 we formally define the fair division problem and strong EFX-orientability. Section 3 discusses the necessary conditions for strong EFX-orientability and obtains the upper bound of 3 on the chromatic number. In the process it completely characterizes 0-1 strong EFX-orientability. Section 4 discusses the sufficiency conditions through bipartiteness or near-bipartiteness of the graph. Section 5 discusses structures in certain 3-chromatic graphs that prevent strong EFX-orientability, leaving 3-chromatic graphs as an ambiguous case. Finally, we conclude with a brief discussion in Section 6.
@@ 2. Preliminaries
A discrete fair division instance is given by $([n],M,\mathcal{F})$, where $[n]$ is the set of agents, $M$ is the set of goods, and =$\mathcal{F}=\{f_{1},f_{2}...f_{n}\}$ is the set of valuation functions, one for each agent. For each $i\in[n]$, $f_{i}:2^{M}\rightarrow\mathbb{R}_{\geq 0}$ represents agent $i$’s value over the bundles (sets) of goods. Thus, $f_{i}$ is non-negative and monotone, i.e., for $A,B\subseteq M$, $A\subseteq B$ implies $0\leq f_{i}(A)\leq f_{i}(B)$. A valuation function $f_{i}$ is said to be additive if for every subset $X=\{m_{1},m_{2},...m_{k}\}\subseteq M$, $f_{i}(X)=f_{i}(\{m_{1}\})+f_{i}(\{m_{2}\})+...+f_{i}(\{m_{k}\})$, and a valuation is 0-1 additive if it is additive and $f_{i}(\{m\})\in\{0,1\}$ for all $m\in M$.
An allocation $X=(X_{1},X_{2}...X_{n})$ is a partition of $M$ where agent $i$ is assigned the bundle $X_{i}$. An allocation is said to be envy-free (EF), if $f_{i}(X_{i})\geq f_{i}(X_{j})$ for all $i,j\in[n]$. Since such an allocation may not exist when goods are indivisible, we consider the following relaxation:
@@ 3.1. Characterization of 0-1 Strongly EFX-Orientable Graphs
A graph $G$ is 0-1 strongly EFX-orientable if and only if, for every subgraph $H\subseteq G$ such that $H$ is a forest consisting of trees $T_{1},T_{2}...T_{k}$, for every $1\leq i\leq k$ there exists $x_{i}\in T_{i}$ such that $\bigcup_{i=1}^{k}N_{H}(x_{i})$ forms an independent set on $G$, where $N_{H}(x)$ denotes the neighbors of $x$ in $H$ (Figure 1 demonstrates this condition for a $K_{2,4}$ graph).
@@
(if direction) Suppose an adversary gives us a 0-1 additive valuation assignment a graph which satisfies these conditions. For all the edges with asymmetric valuation (where the two endpoints do not value the edge equally), orient them towards the vertex that values the edge. Call the vertices that receive the asymmetrically valued edges special vertices. Let $H^{\prime}$ be the subgraph consisting of edges valued at 1 for both endpoints. For each component of $H^{\prime}$ that contains a special vertex, create a spanning tree with the special vertex as the root, and give each vertex the edge from its parent. For a component $C\subseteq H^{\prime}$ that contains a cycle but has no special vertex, we remove one edge $uv\in C$ where $uv$ is in a cycle, and construct a spanning tree through $C-uv$ with $v$ as the root. Assign each vertex in $C$ the edge from its parent, and assign $uv$ to $v$. Orient the rest of the edges in these components arbitrarily. Since all vertices in such components receive an edge they value, they do not envy anyone else, and are hence not envied by any vertex.
@@
(only if direction) Suppose there exists a forest $H$ such that for its trees $T_{1},T_{2}...T_{k}$, for every collection of $x_{1},x_{2}...x_{k}$ such that $x_{i}\in T_{i}$, their neighborhoods in $H$ do not form an independent set on $G$. Set all the edges in $H$ to have weight 1 for both vertices, and set the rest of the edges to have weight 0 for both vertices. Note that any orientation on a tree must have a source, so for any orientation in $G$, we can choose a collection of sources in $s_{1},s_{2}...s_{k}$ in $H$ such that $s_{i}\in T_{i}$ for $1\leq i\leq k$. Since each $s_{i}$ receives nothing of value and has all its valued edges go to its neighbors, all vertices in $N_{H}(s_{i})$ are envied by $s_{i}$, so all vertices in $\bigcup_{i=1}^{k}N_{H}(s_{i})$ must be envied. However, since $\bigcup_{i=1}^{k}N_{H}(s_{i})$ does not form an independent set on $G$, then an edge of weight 0 must be between two vertices in that set. Assigning that edge will break EFX since both its endpoints are envied due to the orientation on $H$, hence such a graph is not 0-1 strongly EFX-orientable. ∎
@@
Let $G$ be a 0-1 strongly EFX orientable graph. For any matching $M$ on $G$, the subgraph induced by the vertices of $M$ has an independent set of size $|M|$.
@@
Let $M$ consist of edges $e_{1},e_{2},....e_{k}$. Each edge in the matching can be considered a tree, and the edges form a forest. Hence, by Lemma 3.1, for each edge $e_{i}$ in $M$, we can select $v_{i}\in e_{i}$ such that $\bigcup_{i=1}^{k}N_{M}(v_{i})$ is an independent set. Since each vertex in $M$ has exactly one other neighbor also in $M$, then $\bigcup_{i=1}^{k}N_{M}(v_{i})$ consists of $|M|$ vertices, and hence is an independent set of size $|M|$. ∎
Our next lemma gives us a method to generate more graphs that violate the conditions of the previous corollary by using subdivisons. The subdivision of an edge $uv\in E(G)$ is defined as deleting the edge $uv$, adding a new vertex $w$, and adding the edges $uw$ and $wv$ in $G$, effectively placing a vertex in the middle of an edge and dividing it into two edges. A graph $G^{\prime}$ is said to be a subdivision of $G$ if $G^{\prime}$ can be obtained through an iteration of edge subdivisions starting from $G$.
@@
Let $G$ be a graph that violates the matching condition in Corollary 3.2. Let $G^{\prime}$ be the graph obtained by subdividing any edge of $G$ twice. Then $G^{\prime}$ violates the matching condition of Corollary 3.2 as well.
@@
Let the subdivided edge on $G^{\prime}$ consist of $u,x_{1},x_{2},v$, where $uv$ is the original edge in $G$ and $x_{1},x_{2}$ are the added vertices. Suppose that for any matching in $G^{\prime}$, there exists an independent set of the same size on the graph induced by the vertices of that matching. Take a matching $M$ on $G$. We split this problem into two cases:
@@
We claim that $K_{4}$ violates the matching condition in Corollary 3.2. Indeed, there is no independent set of size 2 on a $K_{4}$, while it’s possible to have a matching of size 2 in a $K_{4}$. By Lemma 3.3, a totally odd subdivision of a graph that violates the matching condition will violate the matching condition itself as well. Hence, 0-1 strongly EFX-orientable graphs cannot contain a totally odd $K_{4}$ subdivision, so $\chi(G)\leq 3$ by Lemma 3.4. ∎
@@ 4. Strong EFX-Orientability: Sufficient Condition via Bipartiteness
Note that $B$ must be an independent set in $G$, since $uv$ has one end in $A$, and all other edges go from $A$ to $B$, so no envy exists between vertices in $B$. No vertex in $A$ envies a vertex in $B$, since they got to take an edge before $B$, and the remaining edges incident to a vertex in $A$ were assigned to different neighbors of that vertex. While there might be envy towards vertices in $A$, every vertex in $A$ was assigned only one edge, making this orientation EFX. ∎
@@ 5. General Valuations: Forbidden Structures
Finally, the graph of two triangles connected by a path of length 2 has a forest selected that consists of three trees, $T_{1}=v_{1}v_{3}v_{2}$, $T_{2}=v_{4}v_{5}$, and $T_{3}=v_{6}v_{7}$. Similar to the previous case, to select $x_{1},x_{2},x_{3}$ such that their neighborhoods in the forest form an independent set, $x_{1}$ cannot equal $v_{3}$ and must be either $v_{1}$ or $v_{2}$. This means that $x_{2}$ cannot equal $v_{4}$, since $v_{3}$ and $v_{4}$ are adjacent, so $x_{2}=v_{5}$. But both $v_{6}$ and $v_{7}$ are adjacent to $v_{5}$, so selecting $x_{3}$ is impossible, meaning that this graph is not 0-1 strongly EFX orientable as well. ∎
@@
Let the first triangle be $v_{1}v_{2}v_{3}$, and the second triangle be $v_{2}v_{3}v_{4}$. Let $v_{2}v_{3}$ have weight 0.5 for both endpoints, $v_{1}v_{2}$ and $v_{3}v_{4}$ have weight 1 for both endpoints, and the rest of the edges have weight 0 for both endpoints. All valuations are additive. Due to the symmetry of the graph, without loss of generality, we can assume that $v_{2}v_{3}$ is directed towards $v_{2}$. For the orientation to be EFX, $v_{2}$ cannot receive $v_{1}v_{2}$, as $v_{1}$ will envy $v_{2}$ even though $v_{2}$ has already received $v_{2}v_{3}$. Hence $v_{1}$ must receive $v_{1}v_{2}$, but $v_{1}$ will be envied by $v_{2}$, so $v_{1}v_{3}$ must be directed towards $v_{3}$. Since $v_{3}$ already has an edge, for similar reasons, $v_{3}$ cannot receive $v_{3}v_{4}$, so $v_{3}v_{4}$ is directed towards $v_{4}$. Note that $v_{2}$ and $v_{4}$ are both envied by $v_{3}$, but there is another edge between $v_{2}$ and $v_{4}$ that needs to be assigned, making an EFX-orientation impossible.
@@
Note while subdividing $e$, we created edges $ux_{1}$ and $vx_{2}$ which also have zero value, meaning that we can repeat this process again by subdividing any one of those edges twice, hence replacing $e$ with an odd path of edges alternating between zero and one values. The new graph will not have an EFX orientation with this valuation, and hence is not strongly EFX-orientable, as desired. ∎
@@
Let $G$ be a bipartite graph with $|V(G)|\geq 4$ that remains connected after the deletion of any vertex. Suppose that an edge $e$ is added, which connects two vertices of the same partite set. Then $G+e$ is not strongly EFX-orientable.